# Two-Time Relativistic Bohmian Model of Quantum Mechanics

### v10

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The hypothesis that matter is capable of motion with respect to an additional independent time (thus resulting instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking antimatter. The model also predicts a relativistic correction of the uncertainty principle.
@@
Keywords: Quantum Mechanics Foundations; de Broglie-Bohm Theory; Zitterbewegung; Uncertainty principle verification; Extra dimensions, Atomic orbitals, Spin, Definition of time in Quantum Mechanics.
@@ II TTBM definition and properties
Two-time physics has been considered in various contexts. An additional temporal parameter is considered, e.g., in [23, 24, 25, 26, 27, 28, 29] and, with a different approach, in [30, 31, 32, 33, 34, 35] in order to more accurately explain different quantum properties or obtain supersymmetries. But in all these theories, the expectation value of the new time is not independent from the old one.
@@
with ${\rm d}X=v_{c_{x}}{\rm d}t+v_{i_{x}}{\rm d}\tau$. If we suppose that $v_{c_{x}}$ and $v_{i_{x}}$ depend, respectively, only on $t$ and $\tau$, we get
@@
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts as the spreading of the electrons in their atomic orbitals and the quantum diffractions. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [2, 3]. The puzzling aspects of Quantum Mechanics are rooted in ubiquity and the consideration of an extra time is capable of explaining it. The theory starts from the wavefunction
@@
with $R>0$ and $S$ real functions, admitting the validity of the Klein-Gordon equation. We maintain that the meaning of $\psi$, and in particular of $R$, is the same as in the standard case. Beginning with the free case $(V=0)$, by taking the gradient and then the divergence of eq. (3) and, separately, differentiating it twice with respect to $t$, one can obtain [4]
@@
In presence of a scalar potential $V\neq 0$, one has: $H=m\gamma_{c}c^{2}+V_{Q}+V$ and these equations continue to hold with HVH-V replacing $H_{free}$, by using the generalized Klein-Gordon equation. Eq. (5) is nothing other than the well-known continuity equation that can also be obtained by applying the Klein-Gordon equation to $\psi$ and $\psi^{*}$ and which fails to represent the conservation of a probability density (see, e.g. [36]). In fact, we must expressly request the validity of the continuity equation:
@@
which imposes the conservation of the probability of finding the particle in a unit volume. Eq. (6) also holds in the relativistic case, because it is independent of the relativistic mass of the particle and could depend only on its dimensional Lorentzian contraction; which is neglected when treating the particle as a material point.
The equation simplifies when the variability of $\vec{v}_{c}$ from space is not considered:
@@
Contrary to what is often implied, the Klein-Gordon equation is not in conflict with conservation of probability. In fact, eqs. (5) and (6) are not incompatible, but together they give (see Appendix)
which relates the temporal variation of the Quantum Potential to that of the relativistic mass.
@@
After these outcomes, we can identify the expression of the two-time action in eq. (3), that is $S=S_{t}+S_{\tau}$, where
@@ II.1 Some fundamental properties
    $\begin{split}\hskip 25.6073ptV_{Q}=\\ &\hskip-25.6073pt-m\gamma_{c}c^{2}+m\gamma_{c}c{}^{2}\sqrt{1+\frac{\hbar^{2}}{m^{2}\gamma_{c}^{2}c^{4}R}(\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\nabla{}^{2}R)},\end{split}$ (15)
@@
    $\begin{split}-\frac{\hbar{}^{2}}{2m}\frac{\nabla^{2}\psi}{\psi}=H-V-V_{Q}-\frac{\hbar{}^{2}}{2m}\frac{\nabla^{2}R}{R}+\\ +{\rm i}\frac{\hbar}{2mc^{2}R^{2}}\frac{\partial(R^{2}(m\gamma_{c}c^{2}+V_{Q}))}{\partial t},\end{split}$ (16)
@@ III Free particle
Using eqs. (4) and (7), after three integrations and approximating for $R\rightarrow R_{M}$, it is found [4]
@@
By imposing $v_{i_{s_{MAX}}}=c$, we obtain $\Delta{\mathsf{S}}=\frac{\lambda}{f\sqrt{2(f+2)}}$, which generalizes the standard Zitterbewegung if f is of the order of unity; however, here this motion arises by self-interaction with its own wavefunction, without the need to evoke the antiparticle. According to point 1 of the theory, the amplitude $\Delta{\mathsf{S}}$ represents the minimum positional uncertainty in $s$ of the particle: only by means of a wave packet description is it possible to take into account larger positional uncertainties and smaller momentum uncertainties. The uncertainty principle is therefore
@@
Finally, let’s discuss an important assumption of the theory: that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid, because $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise just for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). Later we will explicitly find the oscillatory $\tau$-motions of an electron in a circular atomic orbit.
@@ IV TTBM empiricism
TTBM predicts two new effects beyond standard theory. The anisotropy of the uncertainty principle has already been discussed and could be detected by an angular test in high-energy accelerators. The other effect is due to the finiteness of the volume of space in which the Feynman integrals must be calculated. While in standard theory this volume is unbounded, in TTBM it is represented by the positional uncertainty volume swept by the particle during its motion [^3] It is possible that the introduction of the new time parameter and the finiteness of the volume makes the path integral a properly mathematical object, unlike Feynman’s original formulation..
However, both effects (especially the second) are normally negligible and therefore the experimental predictions are practically coincident with the standard ones. Under this assumption, we will describe all the possibilities that can arise regarding the motion of a particle in a vacuum: evolution with self-interactions, interference (e.g. triggered by wall slits) and collapse after a detection (fig. 1).
@@
At a fixed instant, $\Delta{\mathsf{S}}$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-oscillations, is spread: relative to time $t$, i.e. empirically, the particle must be considered as simultaneously present in every point of this volume. The infinite ”self-particles” interact in $\tau$-time with each other in all possible physical ways: each kind of force-mediating particles (photons, vector bosons, etc.) oscillates itself in $\tau$-time according to the same laws found. Mutual exchanges of such particles correspond to what in the standard point of view is described as emission and reabsorption of virtualvirtual particles, obtaining excellent experimental verifications. In addition, we have self-waves $\tau$-interferences, exactly as predicted by the standard interpretation, according to the Feynman description. TTBM just specifies that the different histories are actually realized in time $\tau$ and therefore are instantaneous with respect to $t$-time (i.e. with respect to an observer). In particular, the interferences that cause diffraction in a wall slits experiment are due to different values of the $\tau$-action ($S_{\tau}$) in the wavefunctions. Collapses, besides, occur as a consequence of an energetic change of the particle: after an inelastic impact, a subgroup of waves changes its classical action $S_{t}$ and interferes in a particularly destructive way (interference due to different $S_{\tau}$ also contributes) with all the other waves [^4] In Feynman’s formulation, interference is due to different values of the classical action, which is assumed to be variable. There is no specific explanation for collapses. .
@@ V Atomic orbitals
In a small arc of an orbit of radius $\bar{r}$, eq. (7), written for $R$, is approximated by
$\frac{\partial R}{\partial t}+\frac{v_{c}}{\bar{r}}\dot{R}=0,$ (25)
where a dot denotes $\theta$-derivative and $\frac{{v}_{c}}{\bar{r}}\equiv{\omega}_{c}$ is the angular velocity. Derivating further, we get
@@
$\begin{split}(\mu\gamma_{c}c^{2}+\frac{{k}}{R^{2}})^{2}-\mu^{2}\gamma_{c}^{2}c^{4}=\\ \frac{\hbar{}^{2}}{R}({\omega}^{2}_{c}\ddot{R}-c^{2}(\frac{2R^{\prime}}{r}+R^{\prime\prime}+\frac{\dot{R}\,\rm{cos}\,\theta}{r^{2}\rm{sin}\,\theta}+\frac{\ddot{R}}{r^{2}})),\end{split}$ (27)
@@ V.1 Tangential $\tau$-oscillation
For a small pure tangential displacement, we can neglect $R^{\prime}$ and $R^{\prime\prime}$ and replace $r$, $\rm{cos}\,\theta$ and $\rm{sin}\,\theta$ with $\bar{r}$, $1$ and $\theta$ into eq. (27), obtaining
$\frac{{k}^{2}}{R^{4}}+\frac{2\mu\gamma_{c}c^{2}\,k}{R^{2}}=\frac{\hbar{}^{2}}{R}({\omega}^{2}_{c}-\frac{c^{2}}{\bar{r}^{2}})\ddot{R}-\frac{\hbar{}^{2}}{R}(\frac{c^{2}\dot{R}}{\bar{r}^{2}\,\theta}).$ (28)
@@
$\frac{\hbar^{2}}{\gamma^{2}_{c}\,\bar{r}^{2}}\,\theta\,\ddot{R}+\frac{\hbar^{2}}{\bar{r}^{2}}\dot{R}+\frac{k^{2}}{c^{2}}\,\frac{\theta}{R^{3}}+2\mu\gamma_{c}k\,\frac{\theta}{R}=0.$ (29)
@@
Similarly to what was done for the free case, we set ${k}=\mu\gamma_{c}c^{2}R_{M}^{2}\bar{f}$, where $R_{M}$ is the maximum value of $R$ and $\bar{f}$ an adimensional positive arbitrary constant. Of course, not only $\bar{f}\rightarrow f$ for $\bar{r}\rightarrow\infty$, but more generally $\bar{f}$ and $f$ always have to be comparable, since in any orbital the Coulomb potential is small compared to $\mu\gamma_{c}c^{2}$. Setting $R(0)=R_{M}$, eq. (30) becomes
@@ V.2 Radial $\tau$-oscillation
Imposing ${k}=\mu\gamma_{c}c^{2}R_{M}^{2}\bar{f}$, $R(\bar{r})=R_{M}$ and substituting into the eq. (34), one obtains
@@ IX Final remarks
The proposed theory arises as a result of taking literally Feynman’s sum over history interpretation. The postulation of a new time parameter is a much simpler expedient, from a mathematical point of view, than the criterion of the Bohmian theories mentioned in the introduction, which predict a convoluted space-time and violate the spirit of Special Relativity. The theory is able to satisfactorily explain the general empiricism of Quantum Mechanics, leading interference, collapses and self-interactions (suspiciously called virtualvirtual in the standard view) to a deterministic explanation. Furthermore, TTBM provides testable effects beyond the standard viewpoint and exciting development prospects involve explanation of spin and dark matter.
@@
$\begin{split}m\gamma_{c}c^{2}\frac{\partial R^{2}}{\partial t}+R^{2}\frac{\partial(m\gamma_{c}c^{2})}{\partial t}-\frac{\partial(R^{2}m\gamma_{c}c^{2})}{\partial t}-\\ -R^{2}\frac{{\rm d}(m\gamma_{c}c^{2})}{{\rm d}t}-\frac{\partial(R^{2}V_{Q})}{\partial t}=0,\end{split}$ (41)

### v11

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The hypothesis that matter is capable of motion with respect to an additional independent time (thus resulting instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking virtual antimatter. The model also predicts a relativistic correction of the uncertainty principle.
@@
Keywords: Quantum Mechanics Foundations; de Broglie-Bohm Theory; Zitterbewegung; Uncertainty principle verification; Extra dimensions; Atomic orbitals; Spin; Definition of time in Quantum Mechanics.
@@ II TTBM definition and properties
Two-time physics has been considered in various contexts. An additional temporal parameter is considered, e.g., in [23, 24, 25, 26, 27, 28, 29] and, with a different approach, in [30, 31, 32, 33, 34, 35] in order to more accurately explain different quantum properties or obtain supersymmetries.
@@
with ${\rm d}X=v_{c_{x}}{\rm d}t+v_{i_{x}}{\rm d}\tau$. If we suppose that $v_{c_{x}}$ and $v_{i_{x}}$ depend, respectively, only on $t$ and $\tau$ [^1] If this were not the case, classical dynamics would have long since suggested the existence of $\tau$ time., we get
@@
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts as the spreading of the electrons in their atomic orbitals and the quantum diffractions. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [2, 3]. The puzzling aspects of Quantum Mechanics are rooted in ubiquity and the consideration of an extra time is capable of explaining it.
The theory starts from the wavefunction
@@
with $R>0$ and $S$ real functions, admitting the validity of the Klein-Gordon equation. Beginning with the free case $(V=0)$, by taking the gradient and then the divergence of eq. (3) and, separately, differentiating it twice with respect to $t$, one can obtain [4]
@@
In presence of a scalar potential $V\neq 0$, one has: $H=m\gamma_{c}c^{2}+V_{Q}+V$ and these equations continue to hold with HVH-V replacing $H_{free}$, by using the generalized Klein-Gordon equation.
Eq. (5) is nothing other than the well-known continuity equation that can also be obtained by applying the Klein-Gordon equation to $\psi$ and $\psi^{*}$ and which fails to represent the conservation of a probability density (see, e.g. [36]). However, in TTBM the non-relativistic meaning of $R^{2}$ as volumetric probability density can be recovered. For that, it is enough to expressly request the validity of the continuity equation:
@@
which imposes the conservation of the probability of finding the particle in a unit volume. This equation must also be extended to the relativistic case, because it is independent of the relativistic mass of the particle and could depend only on its dimensional Lorentzian contraction; which is neglected when treating the particle as a material point.
Now, eqs. (5) and (6) give (see Appendix)
which relates the temporal variation of the Quantum Potential to that of the relativistic mass. This equation does not give rise to any contradiction if $R$ (and therefore also $V_{Q}$) depends on $t$ time, as we have assumed. Consequently, in TTBM, the Klein-Gordon equation is compatible with the Schrödingerian meaning of $R^{2}$ and its conservation, contrary to the standard model.
Eq. (6) simplifies when the variability of $\vec{v}_{c}$ from space is not considered:
@@
After these outcomes, we can identify the expression of the two-time Hamilton’s action in eq. (3), that is $S=S_{t}+S_{\tau}$, where
@@ II.1 Some fundamental properties
    $split\hskip 25.6073ptV_{Q}=\\ &\hskip-25.6073pt-m\gamma_{c}c^{2}+m\gamma_{c}c{}^{2}\sqrt{1+\frac{\hbar^{2}}{m^{2}\gamma_{c}^{2}c^{4}R}(\frac{\partial^{2}R}{\partial t^{2}}-c^{2}\nabla{}^{2}R)},$ (15)
@@
[…]
@@ III Free particle
Using eqs. (4) and (7), after two integrations and approximating for $R\rightarrow R_{M}$, it is found [4]
@@
By imposing $v_{i_{s_{MAX}}}=c$, we obtain $\Delta{\mathsf{S}}=\frac{\lambda}{f\sqrt{2(f+2)}}$, which generalizes the standard Zitterbewegung if f is of the order of unity; however, here this motion arises by self-interaction with its own wavefunction, without the need to evoke the virtual antiparticle. According to point 1 of the theory, the amplitude $\Delta{\mathsf{S}}$ represents the minimum positional uncertainty in $s$ of the particle: only by means of a wave packet description is it possible to take into account larger positional uncertainties and smaller momentum uncertainties. The uncertainty principle is therefore
@@
It is admitted by the theory that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid, because $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise just for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). Later we will explicitly find the oscillatory $\tau$-motions of an electron in a circular atomic orbit.
@@ IV TTBM empiricism
TTBM predicts new effects beyond standard theory. The main ones are due to the aforementioned relativistic differences in intrinsic velocities. The anisotropy of the uncertainty principle has already been discussed and could be detected by an angular test in high-energy accelerators. Another relativistic effect related to this anisotropy should be an additional directional interference (see below). Also, there is the effect due to the finiteness of the volume of space in which the Feynman integrals must be calculated: while in standard theory this volume is unbounded, in TTBM it is represented by the positional uncertainty volume swept by the particle during its motion [^4] It is also likely that the introduction of the new time parameter and the finiteness of the volume makes the path integral a properly mathematical object, unlike Feynman’s original formulation..
However, these effects (especially the last one) are normally negligible and therefore the experimental predictions are practically coincident with the standard ones. Under this assumption, we will describe all the possibilities that can arise regarding the motion of a particle in a vacuum: evolution with self-interactions, interference (e.g. triggered by wall slits) and collapse after a detection (fig. 1).
@@
At a fixed instant, $\Delta{\mathsf{S}}$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-oscillations, is spread: relative to time $t$, i.e. empirically, the particle must be considered as simultaneously present in every point of this volume. The infinite ”self-particles” interact in $\tau$-time with each other in all possible physical ways: each kind of force-mediating particles (photons, vector bosons, etc.) oscillates itself in $\tau$-time according to the same laws found. Mutual exchanges of such particles correspond to what in the standard point of view is described as emission and reabsorption of virtualvirtual particles, obtaining excellent experimental verifications. In addition, we have self-waves $\tau$-interferences, exactly as predicted by the standard interpretation, according to the Feynman description. TTBM just specifies that the different histories are actually realized in time $\tau$ and therefore are instantaneous with respect to $t$-time (i.e. with respect to an observer). In particular, the interferences that cause diffraction in a wall slits experiment are due to different values of the $\tau$-action ($S_{\tau}$) in the wavefunctions, which, neglecting the relativistic differences in intrinsic velocities, are due to the different lengths of the paths [^5] At relativistic speeds, non-standard directional interference should also appear, the characteristics of which we will not delve into here..
Collapses, besides, occur as a consequence of an energetic change of the particle: after an inelastic impact, a subgroup of waves changes its classical action $S_{t}$. As it $\tau$-moves away from the collision point, it thins out and, by destructively interfering with the packets with the old classic action, it weakens; until they annihilate each other at a sufficient distance. All this happens only in $\tau$ time and therefore $t$-instantaneously [^6] In Feynman’s formulation, interference is due to different values of the classical action, which is assumed to be variable. There is no specific explanation for collapses. .
@@ V Atomic orbitals
[…]
@@
[…]
@@ V.1 Tangential $\tau$-oscillation
[…]
@@
[…]
@@
[…]
@@ V.2 Radial $\tau$-oscillation
Imposing $\Barf{k}=\mu\gamma_{c}c^{2}R_{M}^{2}\Bar${f}, $R(\bar{r})=R_{M}$ and substituting into the eq. (34), one obtains
@@ IX Final remarks
The proposed theory arises as a result of taking literally Feynman’s sum over history interpretation. The postulation of a new time parameter is a much simpler expedient, from a mathematical point of view, than the criterion of the Bohmian theories mentioned in the introduction, which predict a convoluted space-time and violate the spirit of Special Relativity. The theory is able to satisfactorily explain the general empiricism of Quantum Mechanics, leading interference, collapses and self-interactions (suspiciously called virtualvirtual in the standard view) to a deterministic explanation. Furthermore, TTBM provides testable effects beyond the standard viewpoint and exciting development prospects involve explanation of spin and part of dark matter.
@@
$split m\gamma_{c}c^{2}\frac{\partial R^{2}}{\partial t}+R^{2}\frac{\partial(m\gamma_{c}c^{2})}{\partial t}-\frac{\partial(R^{2}m\gamma_{c}c^{2})}{\partial t}-\\ -R^{2}\frac{{\rm d}(m\gamma_{c}c^{2})}{{\rm d}t}-\frac{\partial(R^{2}V_{Q})}{\partial t}=0,$ (41)